# Global Prompt Refinement with Non-Interfering Attention Masking for One-Shot Federated Learning

### v1

@@ 1 Introduction
Federated learning (FL) enables collaborative modeling across data sources in a privacy-preserving manner. It learns global knowledge by aggregating model weights or gradients instead of moving data to a central entity mcmahan2017communication ; cai2024lgfgad ; 9820618 ; yi2024federated ; yi2023pfedes ; 11091509 . Recently, Federated Prompt Learning (FPL) has emerged to support parameter-efficient fine-tuning in this setting. It not only leverages pre-trained models (e.g., CLIP radford2021learning ) more effectively, but also reduces communication overhead between FL clients and the FL server by transmitting only learnable prompts qi2025federated ; zhao2023fedprompt ; ren2025advances ; fan2025ten . However, heterogeneous data distributions across clients often degrade model aggregation performance and slow down convergence yang2023efficient ; bai2024diprompt ; li2020federated ; guo2025federated ; guo2024sample . This is because significant differences in the update directions across local models can cause conflicts during aggregation pan2024federated ; cui2024harmonizing ; huang2024federated ; dong2022federated_FCIL ; fu2025federated . As a result, existing FPL methods still require multi-round communications between clients and the server to reconcile inconsistencies arising from client-specific training.
To deal with data heterogeneity, early FPL solutions (e.g., PromptFoilo pan2024federated and DP-FPL tran2025privacy ) typically focused on incorporating global prompt information to constrain the update direction of local prompts in order to enhance the consistency and stability of federated optimization pan2024federated ; cui2024harmonizing ; tran2025privacy ; li2024global . However, since global prompt information is typically available only after the second communication round, these methods inherently rely on multi-round communications. This poses a challenge in real-world scenarios with limited communication resources. To this end, one-shot federated learning (OSFL), which requires only a single round of client-server interaction, offers a promising direction for further exploration zhang2022dense ; zeng2024one ; diao2023towards ; yang2024exploring ; cai2024fgad . OSFL significantly reduces the communication overhead by eliminating iterative exchanges, making it attractive for real-world deployments with strict latency or bandwidth constraints yao2025fedmho ; ma2025geometric ; hu2024fedmut . However, directly applying existing OSFL methods to FPL settings is not suitable, as they are typically optimized for seen tasks while neglecting the preservation of generalization to unseen tasks.
@@ 3 Preliminaries
Suppose there are $K$ clients denoted by $\mathcal{C}=\{C_{1},C_{2},\dots,C_{K}\}$, and each client $C_{k}\in\mathcal{C}$ holds a private dataset $\mathcal{D}_{k}$. The goal of traditional FL is to collaboratively train a global model $f$ by minimizing the objective: $\min_{f}\sum_{k=1}^{K}p_{k}\cdot\mathcal{L}(f,\mathcal{D}_{k})$, where $p_{k}=\frac{|\mathcal{D}_{k}|}{\sum_{j=1}^{K}|\mathcal{D}_{j}|}$ is the weight of client $k$, and $\mathcal{L}(\cdot)$ is a task-specific loss function. In FPL, each client $k$ optimizes a set of learnable prompts $\delta_{k}$ using its private dataset $\mathcal{D}_{k}$, while keeping the backbone model $f$ frozen: $\min_{\delta_{k}}\mathcal{L}(f;\delta_{k},\mathcal{D}_{k})$. To support class-level prediction, the prompt for class $c$ is constructed by concatenating the learnable prompt and a fixed class token. Specifically, let $\mathbf{p}_{c}\in\mathbb{R}^{N_{t}\times d_{t}}$ denote the class token for class $c$, and $\delta_{k}\in\mathbb{R}^{N_{p}\times d_{t}}$ denote the learnable prompt from client $k$, where $N_{t}$ and $N_{p}$ are the lengths of the fixed and learnable prompts, $d_{t}$ is the dimension of hidden embedding, respectively. The full prompt can be formed as $P_{c}^{(k)}=[\delta_{k};\mathbf{p}_{c}]$. The text feature for class $c$ is computed by feeding the prompt into a text encoder $\mathcal{T}(\cdot)$: $\mathbf{f}_{t}^{c}=\mathcal{T}(P_{c}^{(k)})$. Given an input image $\mathbf{x}$, the image feature is extracted by the frozen image encoder $f$: $\mathbf{f}_{v}=f(\mathbf{x})$. The similarity score between the image and class-$c$ textual features is calculated as $\rho_{c}=\operatorname{Cosine}(\mathbf{f}_{t}^{c},\mathbf{f}_{v})$, and the prediction probability is computed via softmax over similarities: $p(\hat{y}=c\mid\mathbf{x})=\frac{\exp(\rho_{c}/\tau)}{\sum_{j=1}^{K}\exp(\rho_{j}/\tau)}$. The training objective is the cross-entropy loss between the similarity-based prediction and the ground-truth label: $\mathcal{L}_{\text{ce}}=\ell(\boldsymbol{\rho},\mathbf{e}_{y})$, where $\boldsymbol{\rho}=[\rho_{1},\dots,\rho_{K}]$, $\mathbf{e}_{y}$ is the one-hot label vector, and $\ell$ denotes the cross-entropy loss function. After local training, client $C_{k}$ uploads its learned prompt $\delta_{k}$ to the server. The server then aggregates the local prompts $\{\delta_{k}\}_{k=1}^{K}$ to construct a global prompt $\delta_{g}$, typically by weighted averaging: $\delta_{g}=\sum_{k=1}^{K}p_{k}\cdot\delta_{k}$.
@@ 4.1 The Attention Isolation (AttnIso) Module
Prompt tuning typically focuses on fitting local training tasks. However, data heterogeneity often leads to divergent optimization across clients and weakens its cross-task generalization. To address this, the AttnIso module designs attention masks to selectively block the information flow from learnable prompt tokens to original text tokens (e.g. “a photo of a [CLASS]”). We denote the text attention matrix as $M\in\mathbb{R}^{1+N_{p}+N_{t}}$, where $N_{p}$ is the number of learnable prompts, $N_{t}$ is the length of input textual feature tokens. The additional 1 corresponds to the [EOS] token. In addition, we apply customized attention control for the [EOS] token to ensure proper aggregation of contextual information from both the prompt and the input tokens.
@@ 5.2 Results and Discussion
This section compares GPR-NIAM-enhanced variants (GPR-NIAMP and GPR-NIAMT, derived by integrating GPR-NIAM into PromptFL and TCP, respectively) with nine baselines: Zero-Shot CLIP radford2021learning with hand-crafted text prompt template, e.g., “a photo of a [class]”, PromptFL guo2023PromptFL , RPO lee2023read , FedTPG qiu2024federated , TCP yao2024tcp , PromptFolio pan2024federated , DP-FPL tran2025privacy , FedDISC yang2024exploring , FedELMY wang2024one . For fair comparison, all methods use the same backbone: a frozen CLIP with ViT-B16 radford2021learning . Following prior work, the learnable prompt is a $10\times 512$ parameter matrix, with 10 prompt tokens and a 512-dimensional embedding.
@@ 6 Conclusions and Future Work
Despite the promising results of GPR-NIAM, there remains room for improvement. The communication overhead from prototype transmission may become a bottleneck in large-scale federated systems with limited bandwidth, which we plan to address in future work. And both the multi-label problem kou2025labeldistributionlearningbiased ; ijcai2024p478 and real-world medical imaging tasks feng2023unsupervised ; feng2025neighbor become considerably more challenging in federated settings. In addition, several issues discussed in Section 5.7 remain open and deserve further investigation.
@@ A.2.3 Performance Comparison across Different Local Training Epoch Settings
(a) DTD (Epochs=1) (b) DTD (Epochs=5) (c) DTD (Epochs=20)
Method Base Novel HM Method Base Novel HM Method Base Novel HM
CLIP 53.35 59.05 56.06 CLIP 53.35 59.05 56.06 CLIP 53.35 59.05 56.06
RPO 52.77 59.42 55.90 RPO 53.24 59.78 56.32 RPO 53.93 60.26 56.92
TCP 53.24 53.38 53.31 TCP 57.63 56.40 57.01 TCP 63.07 59.54 61.25
Method Base Novel HM Method Base Novel HM Method Base Novel HM
CLIP 62.39 78.10 69.36 CLIP 62.39 78.10 69.36 CLIP 62.39 78.10 69.36
RPO 69.97 76.21 72.95 RPO 71.72 76.84 74.19 RPO 74.05 77.05 75.52
TCP 75.51 78.10 76.78 TCP 76.09 79.15 77.59 TCP 79.88 78.31 79.09
@@
To evaluate the effectiveness of the proposed GPR-NIAM under different levels of local training effort, this section investigates its performance across varying local training epochs. Specifically, we consider three settings with local training epochs $\in\{1,5,20\}$, under a data heterogeneity level ($\beta=0.5$) and with 10 clients participating in training. As shown in Table 7 and Table 8, both GPR-NIAMP{P} and GPR-NIAMT{T} consistently outperform the baselines across all datasets and training epochs. Notably, GPR-NIAMT{T} delivers the best overall performance in terms of harmonic mean (HM) in most settings, which demonstrates superior generalization across both base and novel classes. On Flowers102 and UCF101, GPR-NIAMT{T} shows strong robustness even when the number of local epochs increases. This highlights its ability to prevent overfitting to local distributions. Across datasets, both variants of GPR-NIAM remain stable and effective even as the local training dynamics change, outperforming methods that are more sensitive to overfitting or class imbalance under prolonged local updates.
A.2.4 Evaluating Method Scalability with Varying Client Numbers
Table 9: Performance comparison between GPR-NIAM and baselines on DTD and Flowers102 datasets under varying number of clients $K\in\{20,30\}$ in the base-to-base/novel generalization setting.
@@ A.2.4 Evaluating Method Scalability with Varying Client Numbers
Table 10: Performance comparison between GPR-NIAM and baselines on StanfordCars and UCF101 datasets under varying number of clients $K\in\{20,30\}$ in the base-to-base/novel generalization setting.
@@
A.2.5 Evaluation of Module Contributions

### v2

@@ 1 Introduction
Federated learning (FL) enables collaborative modeling across data sources in a privacy-preserving manner, which has gained traction across multiple domains qi2024attentive ; cai2024fgad ; guo2024sample ; li2020federated ; fu2025less . It learns global knowledge by aggregating model weights or gradients instead of moving data to a central entity cai2024lgfgad ; 9820618 ; yi2024federated ; yi2023pfedes . Recently, Federated Prompt Learning (FPL) has emerged to support parameter-efficient fine-tuning in this setting. It not only leverages pre-trained models (e.g., CLIP radford2021learning ) more effectively, but also reduces communication overhead between FL clients and the FL server by transmitting only learnable prompts qi2025federated ; zhao2023fedprompt ; ren2025advances ; fan2025ten . However, heterogeneous data distributions across clients often degrade model aggregation performance and slow down convergence yang2023efficient ; guo2025federated . This is because significant differences in the update directions across local models can cause conflicts during aggregation pan2024federated ; huang2024federated ; dong2022federated_FCIL ; fu2025federated ; zhang2024enabling . As a result, existing FPL methods still require multi-round communications between clients and the server to reconcile inconsistencies arising from client-specific training.
To deal with data heterogeneity, early FPL solutions (e.g., PromptFoilo pan2024federated and DP-FPL tran2025privacy ) typically focused on incorporating global prompt information to constrain the update direction of local prompts in order to enhance the consistency and stability of federated optimization pan2024federated ; cui2024harmonizing ; tran2025privacy ; li2024global . However, since global prompt information is typically available only after the second communication round, these methods inherently rely on multi-round communications. This poses a challenge in real-world scenarios with limited communication resources. To this end, one-shot federated learning (OSFL), which requires only a single round of client-server interaction, offers a promising direction for further exploration zhang2022dense ; zeng2024one ; diao2023towards ; yang2024exploring ; 11091509 . OSFL significantly reduces the communication overhead by eliminating iterative exchanges, making it attractive for real-world deployments with strict latency or bandwidth constraints yao2025fedmho ; ma2025geometric ; hu2024fedmut . However, directly applying existing OSFL methods to FPL settings is not suitable, as they are typically optimized for seen tasks while neglecting the preservation of generalization to unseen tasks.
@@ 3 Preliminaries
Suppose there are $K$ clients denoted by $\mathcal{C}=\{C_{1},C_{2},\dots,C_{K}\}$, and each client $C_{k}\in\mathcal{C}$ holds a private dataset $\mathcal{D}_{k}$. The goal of traditional FL is to collaboratively train a global model $f$ by minimizing the objective: $\min_{f}\sum_{k=1}^{K}p_{k}\cdot\mathcal{L}(f,\mathcal{D}_{k})$, where $p_{k}=\frac{|\mathcal{D}_{k}|}{\sum_{j=1}^{K}|\mathcal{D}_{j}|}$ is the weight of client $k$, and $\mathcal{L}(\cdot)$ is a task-specific loss function hu2024fedcross ; qi2025global ; li2025re ; zhangcausality . In FPL, each client $k$ optimizes a set of learnable prompts $\delta_{k}$ using its private dataset $\mathcal{D}_{k}$, while keeping the backbone model $f$ frozen: $\min_{\delta_{k}}\mathcal{L}(f;\delta_{k},\mathcal{D}_{k})$. To support class-level prediction, the prompt for class $c$ is constructed by concatenating the learnable prompt and a fixed class token. Specifically, let $\mathbf{p}_{c}\in\mathbb{R}^{N_{t}\times d_{t}}$ denote the class token for class $c$, and $\delta_{k}\in\mathbb{R}^{N_{p}\times d_{t}}$ denote the learnable prompt from client $k$, where $N_{t}$ and $N_{p}$ are the lengths of the fixed and learnable prompts, $d_{t}$ is the dimension of hidden embedding, respectively. The full prompt can be formed as $P_{c}^{(k)}=[\delta_{k};\mathbf{p}_{c}]$. The text feature for class $c$ is computed by feeding the prompt into a text encoder $\mathcal{T}(\cdot)$: $\mathbf{f}_{t}^{c}=\mathcal{T}(P_{c}^{(k)})$. Given an input image $\mathbf{x}$, the image feature is extracted by the frozen image encoder $f$: $\mathbf{f}_{v}=f(\mathbf{x})$. The similarity score between the image and class-$c$ textual features is calculated as $\rho_{c}=\operatorname{Cosine}(\mathbf{f}_{t}^{c},\mathbf{f}_{v})$, and the prediction probability is computed via softmax over similarities: $p(\hat{y}=c\mid\mathbf{x})=\frac{\exp(\rho_{c}/\tau)}{\sum_{j=1}^{K}\exp(\rho_{j}/\tau)}$. The training objective is the cross-entropy loss between the similarity-based prediction and the ground-truth label: $\mathcal{L}_{\text{ce}}=\ell(\boldsymbol{\rho},\mathbf{e}_{y})$, where $\boldsymbol{\rho}=[\rho_{1},\dots,\rho_{K}]$, $\mathbf{e}_{y}$ is the one-hot label vector, and $\ell$ denotes the cross-entropy loss function. After local training, client $C_{k}$ uploads its learned prompt $\delta_{k}$ to the server. The server then aggregates the local prompts $\{\delta_{k}\}_{k=1}^{K}$ to construct a global prompt $\delta_{g}$, typically by weighted averaging: $\delta_{g}=\sum_{k=1}^{K}p_{k}\cdot\delta_{k}$.
@@ 4.1 The Attention Isolation (AttnIso) Module
Prompt tuning typically focuses on fitting local training tasks. However, data heterogeneity often leads to divergent optimization across clients and weakens its cross-task generalization. To address this, the AttnIso module designs attention masks to selectively block the information flow from learnable prompt tokens to original text tokens (e.g. “a photo of a [CLASS]”). We denote the text attention matrix as $M\in\mathbb{R}^{(1+N_{p}+N_{t})*(1+N_{p}+N_{t})}$, where $N_{p}$ is the number of learnable prompts, $N_{t}$ is the length of input textual feature tokens. The additional 1 corresponds to the [EOS] token. In addition, we apply customized attention control for the [EOS] token to ensure proper aggregation of contextual information from both the prompt and the input tokens.
@@ 5.2 Results and Discussion
This section compares GPR-NIAM-enhanced variants (GPR-NIAMP and GPR-NIAMT, derived by integrating GPR-NIAM into PromptFL and TCP, respectively) with nine baselines: Zero-Shot CLIP radford2021learning with hand-crafted text prompt template, e.g., “a photo of a [class]” li2024knn , PromptFL guo2023PromptFL , RPO lee2023read , FedTPG qiu2024federated , TCP yao2024tcp , PromptFolio pan2024federated , DP-FPL tran2025privacy , FedDISC yang2024exploring , FedELMY wang2024one . For fair comparison, all methods use the same backbone: a frozen CLIP with ViT-B16 radford2021learning . Following prior work, the learnable prompt is a $10\times 512$ parameter matrix, with 10 prompt tokens and a 512-dimensional embedding.
@@ 5.4 Sensitivity Analysis
5.5 Performance Comparison across Different Local Training Epoch Settings
(a) DTD (Epochs=1) (b) DTD (Epochs=5) (c) DTD (Epochs=20)
Method Base Novel HM Method Base Novel HM Method Base Novel HM
CLIP 53.35 59.05 56.06 CLIP 53.35 59.05 56.06 CLIP 53.35 59.05 56.06
RPO 52.77 59.42 55.90 RPO 53.24 59.78 56.32 RPO 53.93 60.26 56.92
TCP 53.24 53.38 53.31 TCP 57.63 56.40 57.01 TCP 63.07 59.54 61.25
Method Base Novel HM Method Base Novel HM Method Base Novel HM
CLIP 62.39 78.10 69.36 CLIP 62.39 78.10 69.36 CLIP 62.39 78.10 69.36
RPO 69.97 76.21 72.95 RPO 71.72 76.84 74.19 RPO 74.05 77.05 75.52
TCP 75.51 78.10 76.78 TCP 76.09 79.15 77.59 TCP 79.88 78.31 79.09
To assess the robustness of the proposed GPR-NIAM under different local training efforts, we evaluate its performance with local epochs $\in{1,5,20}$, a data heterogeneity level of $\beta=0.5$, and 10 participating clients. As shown in Table 4, both GPR-NIAMP and GPR-NIAMT consistently outperform baselines across all datasets and settings. Notably, GPR-NIAMT achieves the highest harmonic mean (HM) in most cases, indicating better generalization across base and novel classes. On the Flowers102 dataset, it maintains strong performance with increasing local epochs, reflecting robustness against overfitting. Overall, both variants remain stable and effective despite changing local training dynamics.
@@ 6 Conclusions and Future Work
Despite the promising results of GPR-NIAM, there remains room for improvement. The communication overhead from prototype transmission may become a bottleneck in large-scale federated systems with limited bandwidth, which we plan to address in future work. And both the multi-label kou2025labeldistributionlearningbiased ; ijcai2024p478 , multi-modal meng2025causal ; zhengwang_tnnls2025 ; wang2023multi ; cheng2023voice problems and real-world medical imaging tasks feng2023unsupervised ; feng2025neighbor ; dang2025deep become considerably more challenging in federated settings. In addition, several issues discussed in Section 5.8 remain open and deserve further investigation.
Acknowledgments
This work is supported in part by the Key Research and Development Program of Shandong Province (Grant No. 2024TSGC0667); the Ministry of Education, Singapore, under its Academic Research Fund Tier 1 (RG101/24); the RIE2025 Industry Alignment Fund – Industry Collaboration Projects (IAF-ICP) (Award I2301E0026), administered by A*STAR, as well as supported by Alibaba Group and NTU Singapore through Alibaba-NTU Global e-Sustainability CorpLab (ANGEL).
@@ A.2.3 Performance Comparison across Different Local Training Epoch Settings
To evaluate the effectiveness of the proposed GPR-NIAM under different levels of local training effort, this section investigates its performance across varying local training epochs. Specifically, we consider three settings with local training epochs $\in\{1,5,20\}$, under a data heterogeneity level ($\beta=0.5$) and with 10 clients participating in training. As shown in Table 4 and Table 8, both GPR-NIAMP and GPR-NIAMT consistently outperform the baselines across all datasets and training epochs. Notably, GPR-NIAMT{T} delivers the best overall performance in terms of harmonic mean (HM) in most settings, which demonstrates superior generalization across both base and novel classes. On Flowers102 and UCF101, GPR-NIAMT{T} shows strong robustness even when the number of local epochs increases. This highlights its ability to prevent overfitting to local distributions. Across datasets, both variants of GPR-NIAM remain stable and effective even as the local training dynamics change, outperforming methods that are more sensitive to overfitting or class imbalance under prolonged local updates.
A.2.4 Sensitivity Analysis Regarding the Number of Prompt Tokens.
Table 9: Comparison of different methods under two settings (5×512 and 20×512) on Caltech101, DTD, and UCF101 datasets. Bold numbers indicate the best performance.
Caltech101 DTD UCF101
Base Novel HM Base Novel HM Base Novel HM
5×512
PromptFL 88.41 93.97 91.11 59.25 52.77 55.83 76.78 69.91 73.18
FedTPG 86.65 93.37 89.88 60.18 56.15 58.10 74.30 70.63 72.41
PromptFolio 89.47 93.46 91.42 61.02 56.28 57.62 74.40 70.65 72.48
GRP-NIAM_P 89.90 93.37 91.60 60.96 55.95 59.12 77.98 70.39 73.99
20×512
PromptFL 88.91 94.31 91.53 62.50 52.53 57.08 76.83 69.12 72.77
FedTPG 88.84 92.86 90.81 58.10 54.83 56.41 76.30 73.41 74.41
PromptFolio 89.97 94.22 92.05 56.13 53.98 55.03 75.72 74.41 75.06
GRP-NIAM_P 90.02 94.28 92.10 61.43 56.23 58.71 77.23 74.58 75.88
We have conducted additional sensitivity experiments regarding the number of prompt tokens. Specifically, we adjusted the prompt length to 5×512 and 20×512. The results show that GPR-NIAM remains robust across different prompt lengths, while consistently maintaining an advantage over all baselines in all case.
A.2.5 Evaluating Method Scalability with Varying Client Numbers
Table 10: Performance comparison between GPR-NIAM and baselines on DTD and Flowers102 datasets under varying number of clients $K\in\{20,30\}$ in the base-to-base/novel generalization.
@@ A.2.5 Evaluating Method Scalability with Varying Client Numbers
Table 11: Performance comparison between GPR-NIAM and baselines on StanfordCars and UCF101 datasets under varying number of clients $K\in\{20,30\}$ in the base-to-base/novel generalization.
@@
A.2.6 Evaluation of Module Contributions
@@
A.2.7 Sensitive Analysis of Textual Prompt Template
Table 13: Prompt template comparison on Caltech101, DTD, and UCF101. Best numbers in bold.
Template Caltech101 DTD UCF101
Base Novel HM Base Novel HM Base Novel HM
A photo of a [CLASS] 91.59 94.22 92.89 68.40 53.50 60.04 78.57 70.28 74.19
A picture of a [CLASS] 91.43 94.34 92.86 67.37 52.83 59.22 78.61 69.54 73.79
An image showing [CLASS] 91.52 93.88 92.68 68.57 54.12 60.49 78.77 70.14 74.20
The model’s performance is not highly sensitive to the choice of textual prompt templates. Following prior works such as CLIP and PromptFL, we use “a photo of a [CLASS]” as the default template. Since our method mainly optimizes the learnable prompt tokens ($T_{P}$) while keeping the text tokens ($T_{T}$) fixed, the initial template primarily serves as a starting point for semantic alignment and has limited impact on the final performance. We also conducted a small-scale experiment comparing different templates (e.g., “a picture of a [CLASS]”, “an image showing [CLASS]”), and the variation in Top-1 accuracy was less than 1%, indicating that the GPR-NIAM framework is robust to the choice of templates.